The XYZ Company has just gathered estimates for making a business analysis of a new product. Variable costs are constant at $5 a unit; additional plants will be obtained at a cost of $28,000 and depreciated over four years; the new product will be charged $14,000 a year for its share of general overhead; the marketing program calls for an annual advertising expenditure of $15,000 on advertising, $20,000 on distribution, and a price of $9. The firm will have to be able to sell how many units to break even.
1
Expert's answer
2012-07-13T07:23:01-0400
Revenue = TFC+TVC+Profit
Break even point – Profit =0
Revenue = TFC+TVC
Quantity* Price= TFC+TVC
Quantity= (TFC+TVC)/Price
Quantity = ($28,000/4 +$14,000+$15,000+$20,000 + $5*Q)/ $9
Quantity =($56,000+ $5*Q )/ $9
$9*Q=$56,000+ $5*Q
Q=$56,000/4
Q=14,000